OpenStudy (anonymous):

2. If (x + yi) / i = ( 7 + 9i ) , where x and y are real, what is the value of (x + yi)(x - yi)?

OpenStudy (anonymous):

Instead you can do the following: (x + yi) / i = ( 7 + 9i ) So bring the i to the right hands side x+iy= (7+9i)(i)=-9+7i So now you know x+iy Take conjugate of x+iy to find x-iy =? Then multiply x+iy and x-iy to find your answer

OpenStudy (anonymous):

*hand

OpenStudy (lgbasallote):

what about (71 + 9i^2)(7i - 9i^2)

OpenStudy (lgbasallote):

7i + 9i^2 *

OpenStudy (anonymous):

@lgbasallote , are you asking me or @kingmana

OpenStudy (lgbasallote):

im suggesting a solution :P

OpenStudy (anonymous):

@lgbasallote , xD

OpenStudy (lgbasallote):

since i^2 = -1 it becomes (7i - 9)(7i + 9) use that difference of two squares thingy then keep on simplifying as long as you se an i^2 change it to a -1

OpenStudy (lgbasallote):

change it to -1*